OpenStudy (anonymous):

A satellite is in circular orbit at a height H above the Earth’s surface. Find its orbital period (I am looking for a formula, not a number!)

OpenStudy (p0sitr0n):

v^2 = G(M +m)/r u find velocity with C =pi *2r fing C of Earth v = deltax/deltat delta x = C T = v * C

OpenStudy (p0sitr0n):

r is R of Earth + H